A shop sold 20% of its stock on Monday,20% of the remainder on Tuesday and 20% of the remainder on wednesday.On Thursday,they had more than 50% of the original stock left. Explain how this is.
1
Expert's answer
2019-05-01T10:19:38-0400

Let x be the volume of the stock.

On Monday 20% was sold: x*20%/100%=0.2x, rest: x-0.2x=0.8x;

On Tuesday, 20% of the rest was sold:  0.8x*0.2=0.16x, rest: 0.8x-0.16x=0.64x

On Wednesday 20% of the rest was sold: 0.64x*0.2=0.128x, rest: 0.64x-0.128x=0.512x

So, the rest part is 51.2% and it is more than 50% of the original stock.